User Tools

Site Tools

Maintenance items
  • tension of dome wheel
  • oil dome rollers
  • oil shutter bearings
  • lube RA drive
  • tighten focuser set screws
  • check oil in dome gear box
  • keep things clean

5/3/2024: Replaced caster wheel that is just south of due east. We have an extra spare caster wheel in the tool box in the dome. Also, an old car jack in that tool box as well. Note, we suspect this is the one that was replaced 6 years ago on 4/30/2018. Also note, this wheel looks slightly different than the others. We should regularly check this well to ensure it has a tight connection.

1/20/2024: ACP runtime error caused TMO to stop. Telescope got lost from its home position. The dials give the true position. So in the SciTech GotoSync tab you enter the true RA=(LST+HA) and DEC, and then hit sync.

1/14/2023: greased RA drive after a couple of episodes of telescope stalling from west to east

12/2023 : swapped webcams since the original tmocam was no longer able to move

late 2022 : switched back SciTech cable, as communication drops were MUCH more frequent with other cable!

5/21/2022: Swapped out to new USB extender for SciTech to see if it helps with losing communications. Installed Tripp-Lite Model U330-05M cable (shorter and USB 3, although I don't think device supports USB 3).

5/13/2022: changed MaxIM DL to connect to telescope (for guiding!). Added oil to dome gear box. Greased RA drive.

2/4/2022 : changed MaxIM DL configuration to NOT connect to focuser. Seems to solve focuser hang up problems!

1/30/2022: adjusted dome wheel tension a little tighter. Oiled dome wheels.

1/29/2022: after seeing repeated dome failures, inspected and found that gear box had lost its seal and all of its oil. Removed gear box, found some brass shavings from gear. Cleaned carefully and added two new seals with the Right Stuff. However, upon hardening, this did not allow gear to rotate, so removed one of the seals and used old gasket instead (new gasket material purchased, but old one was in good shape and allowed smooth motion). Gear box filled with synthetic 75W-90 oil. Cleaned up dome and installed gear box and motor.

1/6/2022: we've continued to have issues with focuser hanging up. Added an USB 3.0 hub and connected both QHY camera and focusser to it. Also, have been getting Blinky mode issues, so lubricated RA drive some more, still with old Moly EB tube

12/2021 : added eyepiece adapter for guide camera

11/30/2021: dome would not rotate. Unbolting motor showed that set screw holding male gear to the motor shaft was loose, so motor shaft was just rotating inside of it. Tightened that screw and rebolted motor.

11/2021 : issues with focuser hanging up. Attempted to clean USB connectors, perhaps lowered incidence of issues, but did not eliminate them.

10/21/2021: installed ACP 8.3 and ACP Scheduler 8.3

10/6/2021 : over several nights, motors went into blinky mode. Could reproduce by going to west and moving back east: had a repeat of previous issues in this area. Lubricated RA gear with Moly EB, and problem went away.

Replaced Starlight Xpress guide camera with Starlight Xpress Ultra Pro and attempted to focus. Need an extender and focuser; camera needs to be pretty far out and is held in with tape! Also noted significant spherical aberration with guide scope, so precise focus position will be challenging to define.

10/2/2021 : swap to FLI Centerline filter wheel.

09/22/2021 : NOC redirects wireless antenna towards large tower, change from 20 MHz to 40 Mhz, claim to go to almost 300 Mbbs.

02/2020: Simultaneous failure of telescope RA drive and dome rotation. Facilities replaced dome motor. Telescope rebalanced and RA gear greased. See wiki:MaintenanceFeb2020 for full details.

07/2019: Hard drive failure on TMO control computer. Replaced drive, installed Windows 10 (see wiki:Installation).

7/25/18: Software update complete. Updated Windows 7 to latest available updates. Updated ASCOM Platform from 6.0 to 6.4. Updated MaximDL from 5.18 to 5.24. Updated ACP from 6.0 hotfix 4 to 8.1.1. Updated Scheduler from 3.0 to 8.1.1. Scheduler observing plans didn't make the transition - recovery in progress but may not be possible to restore all of them.

7/13/18: Installed 6nm bandpass H-alpha filter (courtesy of Arne at AAVSO) in the slot previously used for a clear filter. Undergoing software/driver/Windows update process.

7/3/18: Cleaned up a lot of dead ants. Installed refurbished Boltwood II sensor.

5/4/18: Dome shutter ran out of power and was left open overnight. Solar panel wires to the battery had been pulled out, probably because they caught on something during dome rotation. Wires have been secured. Spectrograph has been sent back to the owner.

4/30/18: Replaced bad caster wheel. Dome rotation seems significantly improved. Oiled the dome slit chain.

4/25/18: Adjusted tension on spring for dome motor drive wheel. Dome slipping problem may be due to a bad load-bearing caster wheel which is located where the dome slips most often. Current hypothesis is that the weight of the slit and associated parts put extra wear on it; it only rolls intermittently. Bad caster wheel was removed and a replacement is on the way. Dome power is turned off and observatory is non-operational until new wheel is installed, which is planned for tomorrow or this weekend. If replacement goes well, we will probably replace all of the load-bearing caster wheels.

4/11/18: Replaced dome motor drive wheel. Replaced solar panel and added thicker plexiglas to protect it. Investigated telescope RA slew issue; marginal improvement from balance adjustment and reducing slew speed. May look into replacing RA motor, as the current one might be undersized.

3/21/18: oiled dome wheels, dome still intermittently slipping in SE. One of the dome roller wheels seems shot. Spent a bunch of time trying to adjust balance and worm mesh to improve slews from west horizon up, with marginal success only with worm mesh backed all the way out; had to move alarm switch to accommodate that position. Removed solar panel that was broken from a thrown rock…

feb-march 2018: lots of failures with telescope stuck in west, at or near limit

early 2018: failure of Boltwood wind sensor, replaced with spare from Gary, sent failed unit back to Cyanogen

winter 2017: failure of maxdome communication, replaced rotation card with spare from Gary, sent failed card back to Cyanogen

10/15/16: came up to diagnose wireless camera, somehow the wireless hub had SSID switched to astro-wireless-n (not tmo-wireless), so camera wasn't connected. Changed camera configuration to use astro-wireless-n. Mounted Boltwood using epoxy and stainless steel cable ties.

late August: swapped USB-serial adaptor for dome: dome communication now seems to function

08/13/16 : went up after summer. Basic clean-up. UPS appears to have failed. Plugged network switch, antenna into one outlet, computer plug strip and wireless switch into another. Telescope seems to come up fine. Neither webcam seems to work: tmocam can be pinged, but web server doesn't respond (note, recognized/remembered later that must use port 88!); tmocam-wireless does not appear on network. Dome communication fails, even though serial port (COM5) is found. After a few attempts, weather station seems to respond (COM7), but left it inside until everything else is operational. Switched from spectrograph to camera; cameras (QSI and SX) seem to connect fine. Focusser seems to work.

Jan-April 2016: Significant work on dome shutter. Contacted OFS and they installed scaffolding to enable getting to upper part of slit. Charlie Park and his team did inspection, designed and installed significant reinforcement for upper brackets. Top system is similar to bottom, with 6 bearings for the two rails (two below, one above). All 6 lower bearings were replaced. On upper end, 3 were replaced, but middle two (lower) and right lower could not be replaced. These bearings can be lubricated after removing plugs on either end. Charlie lubricated all bearings, and supplied a grease gun to enable this for future maintenance; he recommends annually, but the old bearings hadn't been maintained for 50 years, probably! Charlie also adjusted the cable.

12/23/15: Replaced rubber dome drive wheel; old one had lost a LOT of rubber. Used pry bar to get old one off. Old one had two set screws, but new wheel does not have threaded holes. Used washer to hold pin down. Shutter opens and closes, but it is slow, something is not moving smoothly. Upper bracket on dome is noticeably loose.

12/20/15: dome rotation failure and shutter failure

12/20/15: eastward motion from west of meridian was making noise and had put motors into blinky. Backed off the worm meshing with the balance adjustor all of the way, which seemed to significantly improve the performance.

11/15: installed new webcam, wireless router.

log.txt · Last modified: 2024/05/03 16:31 by astroadmin

Donate Powered by PHP Valid HTML5 Valid CSS Driven by DokuWiki